If r 12 =0.86; r 13 = 0.66; and r 23 = 0.72 show
that r 12.3 = 0.744
1
Expert's answer
2020-11-26T15:57:40-0500

"r_{12.3} = \frac{r_{12} – r_{13} \times r_{23}}{\sqrt{1- r_{13}^2 \times \sqrt{1 - r_{23}^2}}}"


"r_{12.3} = \frac{0.86 – 0.66 \times 0.72}{\sqrt{1 - 0.66^2 \times \sqrt{1 - 0.72^2}}} = \frac{0.38}{0.74 \times 0.69} = 0.744"